Write a variable expression to describe the rule for each sequence. Then find the hundredth term.
1. 3, 6, 9, 12, ...
i am really really really really really really confused. please help??

Answer by stanbon(75887)

Write a variable expression to describe the rule for each sequence. Then find the hundredth term.
1. 3, 6, 9, 12, ...
---
An arithmetic sequence with a = 1 and d = 3
---
a(n) = 1 + (n-1)3 where n is the number of the term.
---
a(100) = 1 + 99*3
a(100) = 1 + 297
a(100) = 298
=====================
Cheers,
Stan H
=====================